OpenStudy (livibear1):

What is the radius of a circle with circumference 13 ft

OpenStudy (chrishogan):

r≈2.07ft squared

OpenStudy (visarika23):

Circumference=2(pi)r=13 So r=13/2(pi)

OpenStudy (faiqraees):

@ChrisHogan It is against the policy of OpenStudy to provide direct answer. Please provide complete explanations so the OP has a better understanding of how the specified problem can be solved.

OpenStudy (chrishogan):

Got it.

OpenStudy (ryderfarris4):

2.07

OpenStudy (djmaster):

Circumference = 2*pi*r pi = 3.14 You can see the step-by-step solution at Calorful: http://goo.gl/JpwB1l You can use this free calculator to solve algebra questions. You can use it to save time from asking.